Pure acid is to be added to a 5% acid solution to obtain 95L of 44% solution. What amounts of each should be used? How many liters of 100% pure acid should be used to make the solution?
Answer by jorel1380(3719)

95x.44=41.8
m+n=95
m+.05n=41.8
100m+5n=4180
100m+100n=9500
95n=5320
n=56
m=39
You need 39 liters of pure acid and 56 liters of 5% acid for your mix..